OpenStudy (anonymous):

Please help with Polynomial functions! Medal and fan! The first part of Ray and Kelsey's roller coaster is a curved pattern that can be represented by a polynomial function. Ray and Kelsey are working to graph a third-degree polynomial function that represents the first pattern in the coaster plan. Ray says the third-degree polynomial has 4 intercepts. Kelsey argues the function can have as many as 3 zeros only. Is there a way for the both of them to be correct? Explain your answer.

OpenStudy (anonymous):

Kelsey has a list of possible functions. Pick one of the g(x) functions below and then describe to Kelsey the key features of g(x), including the end behavior, y-intercept, and zeros. g(x) = x3 − x2 − 4x + 4 g(x) = x3 + 2x2 − 9x − 18 g(x) = x3 − 3x2 − 4x + 12 g(x) = x3 + 2x2 − 25x − 50 g(x) = 2x3 + 14x2 − 2x − 14

OpenStudy (anonymous):

Can a third degree polynomial have 4 intercepts?

OpenStudy (anonymous):

@dan815 @pooja195 please?

OpenStudy (anonymous):

Did you ever get it?